Please help! I'm not sure if you are suppose to divide by the 50 percent or what/
In 1993, 50% of the 96.4 million U.S. households has answering machines. How many households has answering machines?

Answer by Earlsdon(6294)

Hint!
When you see 50% "of" something, the "of" means to multiply, so...
50% of 96.4 million means:
(50%)*(96.4)million. = (0.5)*(96.4)million = 48.2 million.
48.2 million households have answering machines.